# Basic syntax from speech: Spontaneous concatenation in unsupervised deep neural networks

### v2

@@ Abstract
Computational models of syntax are predominantly text-based. Here we propose that the most basic syntactic operations can be modeled directly from raw speech in a fully unsupervised way. We focus on one of the most ubiquitous and elementary properties of syntax—concatenation. We introduce spontaneous concatenation: a phenomenon where convolutional neural networks (CNNs) trained on acoustic recordings of individual words start generating outputs with two or even three words concatenated without ever accessing data with multiple words in the input. We replicate this finding in several independently trained models with different hyperparameters and training data. Additionally, networks trained on two words learn to embed words into novel unobserved word combinations. To our knowledge, this is a previously unreported property of CNNs trained in the ciwGAN/fiwGAN setting on raw speech and has implications both for our understanding of how these architectures learn as well as for modeling syntax and its evolution from raw acoustic inputs.
@@
language evolution , language acquisiton , concatenation , syntax from phonetics , unsupervised deep learning , generative adversarial networks , deep language learning
journal: Journal
@@ 1 Introduction
As our main model of learning we adopt the ciwGAN/fiwGAN model (Beguš, 2021a) which model language as informative imitation and features several advantages over competing computational model and represents one of the closest approximations of human language learning using deep neural networks. The sounds of human speech are a measurable, physical property, but they also encode abstract linguistic information such as syntactic, phonological, morphological, and semantic properties. It has been shown that the model learns symbolic-like rule-like linguistic representations from raw speech (Beguš, 2021a, b, 2020) which illustrates that discrete mental representational units can emerge from continuous physical properties in deep neural networks. The model features several features of human spoken language learning that other models lack: learning by imitation, learning through the production-perception loop, and communicative intent.
@@
Here, we propose that we can model a shift from single unit system to a concatenative system using ciwGAN. We model how compound signals or concatenated words can arise spontaneously in deep neural networks trained on raw speech in a fully unsupervised manner. We thus also provide evidence suggesting that domain-general learning models (such as CNNs) spontaneously concatenate signals in a learning setting that involves learning by imitation and production-perception learning from raw spoken language data (ciwGAN). The paper thus introduces a model of the first step in the development from holistic non-syntactic signals to concatenated signals that represents the most elementary syntactic processing. We also show that the observed spontaneous concatenation is not an idiosyncratic property of a single model, but emerges in several replicated models as well as in models that features several changes in the architecture.
@@ 2 Methods
To test whether CNNs can spontaneously concatenate lexical items, we conduct two sets of experiments. In the first set of experiments (three one-word experiments), we train the networks on single-word inputs and test the networks for concatenated outputs. In the second experiment, we train the networks on one-word and two-word inputs (the two-word experiment) and withhold a subset of two-word combinations. We then test whether words can be embedded into novel unobserved combinations in the output. Such a design also mimics one-word and two-word stages in language acquisition (Berk and Lillo-Martin, 2012).
@@ 2.1 The model
The ciwGAN/fiwGAN architectures involve three networks (Figure 1 and Table 1): the Generator that takes latent codes $c$ (either one-hot vectors or binary codes) and random latent space variable $z$ ($z\sim\mathcal{U}(-1,1)$) and through five or six upconvolutional layers generates 1.024s or 2.048s audio with 16 kHz sampling rate (16,384 or 32,768 samples). The audio is then fed to the Discriminator, which evaluates realness of the output via the Wasserstein loss (Arjovsky et al., 2017). The unique aspect of the ciwGAN/fiwGAN architecture is a separate Q-network, which is trained to estimate the Generator’s hidden code $c$. In the ciwGAN architecture, the latent code $c$ is a one-hot vector; in the fiwGAN architecture the latent code is a binary code. It has been shown that even if there is a mismatch between the number of lexical items and the number of classes, lexical learning nevertheless occurs in the models (Beguš, 2021a).

### v3

@@ Abstract
Computational models of syntax are predominantly text-based. Here we propose that the most basic first step in the evolution of syntax can be modeled directly from raw speech in a fully unsupervised way. We focus on one of the most ubiquitous and elementary suboperation of syntax—concatenation. We introduce spontaneous concatenation: a phenomenon where convolutional neural networks (CNNs) trained on acoustic recordings of individual words start generating outputs with two or even three words concatenated without ever accessing data with multiple words in the input. We replicate this finding in several independently trained models with different hyperparameters and training data. Additionally, networks trained on two words learn to embed words into novel unobserved word combinations. We also show that the concatenated outputs contain precursors to compositionality. To our knowledge, this is a previously unreported property of CNNs trained in the ciwGAN/fiwGAN setting on raw speech and has implications both for our understanding of how these architectures learn as well as for modeling syntax and its evolution in the brain from raw acoustic inputs. We also propose a potential neural mechanism called disinhibition that outlines a possible neural pathway towards concatenation and compositionality and suggests our modeling is useful for generating testable prediction for biological and artificial neural processing of speech.
@@
language evolution , language acquisition , concatenation , syntax from phonetics , unsupervised deep learning , generative adversarial networks , deep language learning
journal: tba
@@ 1 Introduction
Concatenation in itself is not yet syntax. Many other systems, such as bird song (Berwick et al., 2011; Okanoya, 2013; Youngblood, 2024) or music (Patel, 2007) contain concatenation with no referential meaning. The degree to which these systems are analyzed in terms of syntactic processing is still disputed. In our proposal, meaning is modeled in a highly abstract way with binary codes. This is only an imperfect, but not insurmountable approximation of the actual referential meaning. Concatenating raw phonetic items into multiple-word outputs is thus the very first and most basic precursor to syntax, and not syntax itself. However, without this step, no further syntactic evolution could have been possible. Human syntax is also fully compositional. We show that our model not only spontaneously concatenates, but also does so in a compositional manner for at least a subset of lexical items, which presents an additional step towards the evolution of human syntax.
As our model of learning we adopt ciwGAN and fiwGAN (Beguš, 2021a) which model language as informative imitation, feature several advantages over competing computational model, and represent one of the closest approximations of human language learning using deep neural networks. The sounds of human speech are a measurable, physical property, but they also encode abstract linguistic information such as syntactic, phonological, morphological, and semantic properties. It has been shown that the model learns symbolic-like rule-like linguistic representations from raw speech (Beguš, 2021a, b, 2020, 2022) which illustrates that discrete mental representational units can emerge from continuous physical properties such as sound in deep neural networks. The model features several aspects of human spoken language learning that other models lack: learning by imitation, learning through the production-perception loop, and communicative intent.
@@
Here, we propose that we can model a shift from single unit system to a concatenative and partially compositional system using ciwGAN. We model how compound signals or concatenated words can arise spontaneously in deep neural networks trained on raw speech in a fully unsupervised manner. We thus also provide evidence suggesting that domain-general learning models (such as CNNs) spontaneously concatenate signals in a learning setting that involves informative imitation: learning by imitation and with the production-perception loop from raw spoken language data (ciwGAN). The paper thus introduces a model of the first step in the development from holistic non-syntactic signals to compositionally concatenated signals that represents the precursor to syntactic processing. We also show that the observed spontaneous concatenation is not an idiosyncratic property of a single model, but emerges in several replicated models as well as in models that features several changes in the architecture.
In Section 4.1, we provide a plausible neural explanation for the concatenative behavior called the disinhibition which suggests that such modeling can be useful for generating and simulating predictions for neural processing of speech. We also show that the networks exhibit traces of compositionality (Section 3.6), which represent an additional step toward the evolution of human syntax. In Section 4.2, we formalize the interaction between latent variables and connect it to existing mathematical treatments of the Merge operation.
@@ 2 Methods
To test whether CNNs can spontaneously concatenate lexical items, we conduct two sets of experiments. In the first set of experiments (four one-word experiments), we train the networks on single-word inputs and test the networks for concatenated outputs. In the second experiment, we train the networks on one-word and two-word inputs (the two-word experiment) and withhold a subset of two-word combinations. We then test whether words can be embedded into novel unobserved combinations in the output. Such a design also mimics one-word and two-word stages in language acquisition (Berk and Lillo-Martin, 2012).
@@ 2.1 The model
The ciwGAN/fiwGAN architectures involve three networks (Figure 1 and Table 3): the Generator that takes latent codes $c$ (either one-hot vectors or binary codes) and random latent space variables $z$ ($z\sim\mathcal{U}(-1,1)$) and through five or six upconvolutional layers generates 1.024s or 2.048s audio with 16 kHz sampling rate (16,384 or 32,768 samples). The audio is then fed to the Discriminator, which evaluates realness of the output via the Wasserstein loss (Arjovsky et al., 2017). The unique aspect of the ciwGAN/fiwGAN architecture is a separate Q-network, which is trained to estimate the Generator’s hidden code $c$. In the ciwGAN architecture, the latent code $c$ is a one-hot vector; in the fiwGAN architecture the latent code is a binary code. It has been shown that even if there is a mismatch between the number of lexical items and the number of classes, lexical learning nevertheless occurs in the models (Beguš, 2021a).